# SingVisio: Visual Analytics of Diffusion Model for Singing Voice Conversion

### v1

@@ 8.4 Insight and Inspiration
Additionally, E4 noted that the metric comparison perspective reveals the limitations of existing objective metrics used in voice conversion. For example, in a 1000-step diffusion model, almost all metric curves approach convergence within about the last 30 steps, showing no significant improvement beyond that point. However, from the step comparison perspective, we can see (and hear) a substantial difference in sound quality between the generated results at step 30 and step 0, indicating areas for further enhancement.
@@ 9 Conclusion
In this work, we introduce SingVisio, a visual analysis system designed to interactively explain the diffusion model for singing voice conversion. Specifically, SingVisio visually exhibits the stepwise generation process of diffusion models, illustrating the gradual denoising of the noisy spectrum, ultimately resulting in a clean spectrum that captures the target singer’s timbre. The system also supports pairwise comparisons between different conditions, such as content and melody in source audio, and timbre from the target audio, revealing the impact of these conditions on the diffusion generation process and converted results. Through comprehensive evaluations, SingVisio demonstrates its effectiveness in terms of system design, functionalities, explainability, and usability. It provides diverse users with fresh learning experiences and valuable insights into the diffusion model for singing voice conversion.
Appendix A Objective Questions
Table 1 provides a comprehensive overview of the objective questions corresponding to five distinct analytical tasks outlined in Section 4.2, labeled from O1 to O17. The first row of the table describes the task, the second row provides background information for answering questions under different tasks, and from the third row to the last, there are detailed descriptions of specific questions for each task.
Table 1: Objective questions correspond to five analytical tasks in Section 4.2, denoted as O1 to O17.
Task
T1: Step-wise Diffusion Generation Comparison. T2: Step-wise Metric Comparison. T3: Pair-wise SVC Comparison with Different Target Singers. T4: Pair-wise SVC Comparison with Different Source Singers. T5: Pair-wise SVC Comparison with Different Songs.
Context
In the Step Comparison mode, answer the following questions (O1-O4).
In the Metric Comparison mode, answer the following questions (O5-O8).
In the Target Singer Comparison mode, answer the following questions (O9-O11).
In the Source Singer Comparison mode, when converting the singing voice of source singers 8 and 12 with song 12 to target singer 13, respectively, please answer the following questions (O12-O14).
In the Song Comparison mode, when converting the timbre of the source singer 12 to target singer 13 under the conditions of song 12 and song 13, answer the following questions (O15-O17).
Objective Questions
O1: How does the audio quality of the converted singing voice evolve throughout the diffusion generation process?
O5: What is the relationship between the F0CORR (F0 Correlation) metric and the model’s performance?
O9: Is there a difference in timbre between target singer 8 and target singer 13?
O12: What are the timbre types of source singer 8 and singer 12, respectively?
O15: Are the contents of the two sources (Singer 12 Song 12 and Singer 12 Song 13) consistent with the content of the target (Singer 13)?
O2: What is the trend in the changes of the F0 curve and Mel spectrogram during the diffusion generation process?
O6: What aspect of audio does the FAD (Fréchet Audio Distance) metric evaluate?
O10: What are the fundamental frequency ranges of target singer 8 and target singer 9, respectively?
O13: From the project embedding of Step + Noise + Condition (Last layer) shown in the projection view, what conclusion can be drawn from the motion trajectories of the steps corresponding to two conversion processes?
O16: For the final two converted results at step 0 (Song 12: Singer 12 -> Singer 13 and Song 13: Singer 12 -> Singer 13), into whose singer identity should they be converted?
O3: What’s the F0 range in the final converted results (Step 0) with the condition of source singer 1, song 1, and target singer 5?
O7: What is the trend of the MCD (Mel-cepstral Distortion) curve with the change of time steps during the diffusion generation process?
O11: With source singer as 12, song as 12, and target speakers as 8 and 9, respectively, at diffusion step 30, which result has clearer harmonics (horizontal bright lines) in the Mel spectrogram?
O14: What are the timbre types and fundamental frequency range of the final generated results at diffusion step 0?
O17: For the final two converted results at step 0 (Song 12: Singer 12 -> Singer 13 and Song 13: Singer 12 -> Singer 13), which singing content do they match?
O4: At which step does the result of the converted singing voice become intelligible without noise under the system’s default display?
O8: What does the changing trend of Dembed (Singer Similarity) during the diffusion generation process indicate?
@@ Appendix B Dataset
We use five datasets for our diffusion-based SVC model training: Opencpop Wang et al. (2022c), SVCC training data Huang et al. (2023b), VCTK Yamagishi et al. (2019), OpenSinger Huang et al. (2021), and M4Singer Zhang et al. (2022). In total, these datasets contain 83.1 hours of speech and 87.2 hours of singing data. The mapping between the singer name defined in the dataset and the singer ID displayed in the SingVisio system is listed in Table 2. The mapping between the song name defined in the dataset and song ID in the SingVisio system is listed in Table 3.
@@ Appendix C Metrics Definition
SVCC
30001
Song 1
Hey Jude, don’t make it bad.
30002
Song 2
Take a sad song and make it better.
30003
Song 3
Remember to let her into your heart.
10001
Song 4
Everything is fine.
10030
Song 5
Were you lying all the time?
10120
Song 6
Now, I need
10140
Song 7
Hey, I love you.
30005
Song 15
You know that its fool who plays it cool.
30006
Song 16
Na, na, na, na, na, na, na, na, na, na, na, na, hey, Jude.
30009
Song 17
When they all should let us be.
30016
Song 18
Let it be. Let it be. Let it be.
30022
Song 19
Take my breath away
30019
Song 20
Watching every motion In my foolish lover’s game
M4Singer
Alto-1_美错 _0014
Song 8
美丽的错误往往最接近真实
Bass-1_十年 _0008
Song 9
陪在一个陌生人左右
Soprano-2_同桌的你 _0018
Song 10
谁遇见多愁善感的你
Tenor-5_爱笑的眼睛 _0010
Song 11
这爱的城市虽然拥挤
Alto-7_寂寞沙洲冷 _0000
Song 12
河畔的风放肆拼命的吹, 无端拨弄离人的眼泪
Tenor-6_寂寞沙洲冷 _0002
Song 12
Alto-7_寂寞沙洲冷 _0011
Song 13
当记忆的线缠绕过往支离 破碎, 是慌乱占据了心扉
Tenor-7_寂寞沙洲冷 _0013
Song 13
Tenor-6_寂寞沙洲冷 _0020
Song 13
Bass-1_寂寞沙洲冷 _0021
Song 14

### v2

@@ 7 Expert Study
Additionally, E4 noted that the metric comparison perspective reveals the limitations of existing objective metrics used in SVC. For example, in a 1000-step diffusion model, almost all metric curves approach convergence within about the last 30 steps, showing no significant improvement beyond that point. However, from the step comparison perspective, we can see (and hear) a substantial difference in sound quality between the generated results at step 30 and step 0, indicating areas for further enhancement. This observation suggests that while numerical metrics may indicate stabilization, the perceptual quality of audio continues to improve significantly in the final stages of the diffusion process. E4 emphasized that this discrepancy highlights a critical gap in current evaluation methods, as metrics like MCD, FAD, and F0RMSE may not fully capture the nuanced improvements audible to human listeners. To address this, E4 suggested developing new, perceptually aligned metrics that better reflect auditory differences observed during the final diffusion steps.
8 Evaluation
This section details the evaluation approach and the results of SingVisio in both basic and advanced versions. The evaluation is carried out through structured user studies, designed to assess both objective understanding and subjective experiences of the users.
8.1 User Study Set-up
Participants We recruited 23 participants (P1-P23) from audio, music, and speech processing laboratories. They included beginners new to the field, doctoral students with 1-2 years of experience, and postdoctoral researchers with 4-6 years of experience. This mix of participants ensured a broad perspective on the system’s performance across different user groups. Additionally, their research interests centered on audio, music, and speech processing. While they had limited knowledge of visual analysis, they showed great interest in the SingVisio system as it allowed them to interactively visualize their research content, e.g., audio, Mel spectrograms, generative models.
Questionnaire The questionnaire was designed to capture both objective and subjective aspects of the SingVisio. The questions are designed following ContextWing [56] and also considering the specific features of our own system.
  • 1.
    Objective Questions  In the part of objective questions, to evaluate the effectiveness of both the basic and advanced versions, two sets of questionnaires were designed. These questions are directly related to the five analytical tasks (T1-T5) previously detailed in Section 4.2.
    • 1..1
      Objective Questions (Basic Version) (OB1-OB8) For participants in the basic version, the study was conducted as a comparative analysis, where users were divided into two subgroups. One group engaged with SingVisio, while the other group utilized the traditional tutorial method to learn about diffusion-based SVC models, with the same dataset (audio files, Mel spectrograms with F0 visualized, metric data spreadsheet) of every step provided, to answer the same set of questions. This comparative approach allowed us to directly assess the efficiency of SingVisio in helping beginners grasp concepts compared with conventional learning methods.
    • 1..2
      Objective Questions (Advanced Version) (OA1-OA15) The questionnaire for the advanced version is designed with users with more experience or specialized in audio, music or speech processing, aiming to evaluate the effectiveness of the system in aiding field experts in facilitating a more sophisticated analysis and understanding.
  • 2.
    Common Subjective Questions (S1-S16) Both versions of the questionnaire included a shared set of subjective questions, which are intended to capture the users’ perceptions, satisfaction, and any qualitative feedback regarding their experience. The questions are designed following ContextWing [56] in evaluating the system around four key aspects, including explainability, analysis functionality, design effectiveness, and usability. These aspects were selected based on the recommendations by Rossi et al. [57], ensuring a comprehensive evaluation framework that aligns with established user experience principles. The questions are rated using a 5-point Likert scale, ranging from 1 (strongly disagree) to 5 (strongly agree).
By employing this dual-version structure, our questionnaire not only provides insights into the specific utilities of each version of SingVisio but also allows for a nuanced analysis of its educational impact compared to traditional methods. This methodology supports a robust evaluation of the system’s effectiveness across a spectrum of users, from novices to advanced practitioners.
Procedure We first divided the participants by their experience in the field of audio, music, or speech processing in terms of years ($<$1yr, basic group, >=>=> =1yr, advanced SingVisio group), the basic group is further split evenly and randomly into two sub-groups, basic SingVisio group and tutorial group. Out of the 23 participants, the advanced group consisted of 10 individuals. The basic group included 13 individuals, with 7 in the basic SingVisio group and 6 in the tutorial group. Both advanced and basic SingVisio groups were first oriented with a comprehensive introduction to the SingVisio system, familiarizing the participants with the functions and capabilities of SingVisio. The tutorial group was oriented with a tutorial session to learn about necessary knowledge about diffusion-based SVC models and basic concepts like F0, Mel spectrograms and metric definitions. Following the initial setup, participants proceeded to complete the online questionnaire. Those in the SingVisio groups answered the questions while actively using the SingVisio system, configured as specified in the questionnaire. Conversely, participants in the tutorial group answered the questions using a tutorial handout, whilst having access to the same dataset as the SingVisio groups. This dataset included audio files, Mel spectrograms with visualized F0, and a spreadsheet detailing metric data for each step of the process. After both objective and subjective queries were completed, an optional feedback section was provided for any additional comments or suggestions.
8.2 Results and Analysis
The completion time for the basic tutorial group and basic SingVisio group on the basic version was approximately 94.31 ($\sigma=78.23$) minutes and 48.65 ($\sigma=21.93$) minutes, respectively. Additionally, the completion time for the advanced SingVisio group was about 40.54 ($\sigma=26.62$) minutes. It is noted that the completion time for the user study includes not only the time taken to answer the questionnaire but also the time spent familiarizing with the SingVisio system or tutorial. Additionally, the study was conducted in an uncontrolled environment, where participants used their own computers, resulting in potential distractions that could have affected the completion times. Even though removing the extreme outlier, the completion time of the tutorial group ($\mu=68.13$, $\sigma=50.13$) is greater than that of the other two groups. It indicates that the visual and interactive approach in the SingVisio system is more conducive to completing the questionnaire, thereby resulting in shorter completion times.
The average accuracies for the tutorial group and basic SingVisio group were 71.73% and 82.14%, respectively, and the average accuracy for this group was approximately 91.77%. The results indicate that using the SingVisio system requires significantly less time to complete the user study compared to the traditional tutorial method. Meanwhile, SingVisio effectively aids beginners in understanding diffusion-based SVC more efficiently. In contrast, the traditional tutorial method involves manually finding and comparing audio or Mel spectrograms from thousands of files. SingVisio simplifies this by allowing the dynamic display of audio and Mel spectrograms at specified steps, thereby improving efficiency. Furthermore, the higher accuracy and reduced completion time observed among users of the advanced SingVisio group can be attributed to their professional background in signal processing. With at least two years of experience, these users are better equipped to efficiently navigate the system and effectively extract relevant information, contributing to their overall performance.
Fig. 4: Accuracy of objective questionnaires on the basic version, including tutorial group and basic SingVisio group. The questions designed for the basic version are related to analysis tasks T1 and T2 as described in Section 4.2.
8.2.1 Objective Evaluation for basic version
The accuracy of the objective questions (OB1-OB8) for the basic version is shown in Figure 4, including the tutorial group and SingVisio group. Overall, all questions from the SingVisio group obtained higher accuracy than those from the tutorial group except for a question (OB4). The detailed questions and results are described as follows.
Step-wise Diffusion Generation Comparison (T1). We designed four questions (OB1-OB4) related to the diffusion generation process in SVC for the basic version. The objective results shown in Fig. 4 show that the accuracy of OB1 from both groups were 100%, indicating that the system’s capability for users to learn about the diffusion generation process is on par with the tutorial. The accuracies of OB2 and OB3 from the SingVisio group were higher than those from the Tutorial group, indicating that the SingVisio system allows users to clearly observe the F0 range of the audio (as shown in the annotation in Fig. 2). While the tutorial group achieved slightly higher accuracy than the basic SingVisio group on OB4, further analysis provided insight into this discrepancy. The tutorial group benefited from a t-SNE visualization example with handwritten digit recognition, which clearly demonstrated clustering. In contrast, SingVisio’s t-SNE pattern (as shown in the right bottom part of Fig. 2 ) in the diffusion generation process, while forming clusters, was less apparent. This greater clarity in the tutorial’s visual representation likely led to higher accuracy for this question in the tutorial group.
Step-wise Metric Comparison (T2). For task T2, we designed four questions (OB5-OB8). Comparing the accuracy rates for OB5-OB8, we found that the SingVisio group consistently outperformed the Tutorial group, with the most significant gains in OB6, followed by OB7, OB5, and OB8. OB6 and OB7, which focus on the overall trend of the metric curve (as annotated in Fig 3), showed that SingVisio’s interactive and intuitive display of the complete curve is more effective than the tutorial’s method of using Excel sheets to deduce trends. OB5 and OB8 involve understanding the relationship between metrics and model performance. SingVisio’s helpful tool-tips explaining terms or concepts aid in better understanding, resulting in higher accuracy rates for SingVisio.
Fig. 5: Accuracy of objective questionnaires on advanced version. The questions designed for the advanced version are related to all analysis tasks T1-T5, as described in Section 4.2.
8.2.2 Objective Evaluation for advanced version
The result of the objective evaluation for the advanced version is presented in Fig. 5. The data reveal that the majority of questions were answered with accuracies between 90% and 100%, with only three questions falling below this threshold. This suggests that SingVisio effectively supports researchers in answering queries pertinent to diffusion models and SVC. Considering the proficiency of advanced users in these subjects, the questions designed for this version (T1 and T2) were intentionally made more complex than those in the basic version. These 15 objective questions, designated OA1 to OA15, are described as follows.
Step-wise Diffusion Generation Comparison (T1). Questions OA1-OA3 related to T1 are designed for the advanced version. From the result shown in Fig. 5 of the three questions relating to T1, OA1 achieved 90% accuracy, while both OA2 and OA3 achieved 100% accuracy. This indicates the effectiveness of SingVisio in helping users acquire knowledge about diffusion generation and understand the corresponding influence of the Mel spectrogram and F0 contour.
Fig. 6: Rating scores of subjective questionnaires. A1-A4 represents the four aspects, including explainability, functionality, effectiveness, and usability. S1-S15 denotes 15 subjective questions. The rightmost value shows the mean ± standard deviation for each question.
Step-wise Metric Comparison (T2). Three questions related to T2 are designed for the advanced version (OA4-OA6). The ranking in the accuracy of the three questions related to T2, OA6, OA5 and OA4, were 100%, 90% and 60%, respectively. The score for OA4 was relatively low. After consulting several users, we found that the descriptions ”metrics show improvement” and ”metrics show degradation” referred to whether the metric itself improves or degrades during the diffusion generation process, not whether its value increases or decreases. This misunderstanding led to a lower accuracy rate for OA4. In contrast, OA6 received a 100% accuracy rate, indicating that the provided projection embedding is interpretable and allows users to recognize patterns.
Pair-wise SVC Comparison with Different Target Singers (T3). Regarding T3, there are questions (OA7-OA9). Among these questions, OA7 and OA9 both achieved 90% accuracy, while OA8 achieved 80% accuracy. OA7 pertains to the timbre of the singing voice. SingVisio provides both a visual Mel spectrogram and audible audio. In SingVisio, users can select the specified singer via the control panel and listen to the corresponding audio, allowing flexible and efficient analysis while comparing it with the Mel spectrogram. OA9 involves analyzing the difficulty of converting the same source to different target singers. The 90% accuracy indicates that SingVisio effectively helps users determine which conditions in SVC are easy and which are challenging.
Pair-wise SVC Comparison with Different Source Singers (T4). T4 aims to analyze and understand SVC under different source conditions. For T4, three questions (OA10-OA12) were designed. From the results, we can find that all questions have high accuracy. Specifically, OA11 gets 100% accuracy, and OA10 and OA12 obtain 90% accuracy. OA11 involves analyzing whether two singers have different singing styles. This can be observed from the Mel spectrogram, where the harmonic patterns in density and position are noticeably different, and from the audio, where the differences in singing styles can be heard. OA10 pertains to analyzing the F0 (fundamental frequency) of two singers. This can be determined by observing the red-marked F0 contour in the Mel spectrogram. OA12 involves analyzing the duration and fundamental frequency of the conversion results. This information can also be obtained from both the Mel spectrogram and the audio. These results demonstrate that SingVisio provides an effective and flexible tool that offers both audible and visual insights, enabling users to gain comprehensive information from various perspectives.
Pair-wise SVC Comparison with Different Source Songs (T5). For T5, three questions were designed(OA13-OA15). The accuracy rankings for these questions are OA15, OA13, and OA14, with scores of 100%, 90%, and 70% respectively. OA15 involves comparing the projection embedding patterns of two conversion processes. The projection view shows similar trajectories for the hidden features in both conversions, indicating that SingVisio’s projection view is highly effective for analyzing hidden features in diffusion generation.
OA13 and OA14 pertain to the timbre and singing content of two conversion results. Although these questions should ideally have no errors, user inquiries revealed that users often assume the source in SVC includes both content and melody information. Our data includes the same singer performing different songs, which is why our control panel has separate settings for source singer and song. Users mistakenly assumed the source singer included the target content. For future studies, we will avoid ambiguous terms and clearly explain the conditions and questions.
8.2.3 Subjective Evaluation
We conducted subjective evaluations of four aspects (A1-A4), including explainability, functionality, effectiveness, and usability. Overall, the subjection evaluation comprises 15 subjective questions (S1-S15), each scored on a scale ranging from 1 to 5, i.e., strongly disagree (1), disagree (2), neutral (3), agree (4), and strongly agree (5). The assessment results yield an average score of 4.67 ($\sigma=0.02$). Notably, it achieved the highest score in analysis functionality ($\mu=4.76$, $\sigma=0.13$) and also performed well in effectiveness ($\mu=4.74$, $\sigma=0.0$). Detailed results for each dimension and question are presented in Fig. 6.
Explainablility (A1) As shown in Fig. 6, the subjective assessment results across four dimensions indicate that explainability obtains the high score ($\mu=4.70$, $\sigma=0.06$), demonstrating the effectiveness of SingVisio in interpreting diffusion models and SVC. Among the four questions designed to validate the explainability of SingVisio, S1-S4, S2 scored the highest ($\mu=4.78$, $\sigma=0.52$), indicating the system’s effectiveness in aiding users to understand and explain metrics changing over the diffusion generation process. S1, S3, and S4 all received more than 4.45 scores, further confirming that our system provides a comprehensive understanding and explanation for the diffusion model in the context of the SVC task.
Analysis Functionality (A2) The test results revealed that in the subjective assessment across four dimensions, the score in the analysis functionality dimension is the highest ($\mu=4.76$, $\sigma=0.13$). This dimension’s subjective evaluation is designed to verify the system’s support for analysis across tasks T1-T5 and includes four specific questions, S5-S8. Among these questions, S6, S7, and S8 scored the same highest score ($\mu=4.83$), with all users agreeing or strongly agreeing that SingVisio supports analysis and comparison of generated results at different diffusion steps (T1) and the analysis of evaluation metrics (T2). S8, designed for the analysis tasks T3-T5, received agree and strongly agree from all but one neutral user, indicating that SingVisio effectively supports analysis for T3-T5.
Visual Design Effectiveness (A3) To validate the effectiveness of our system design, we formulate four questions (S9-S12). All four questions scored about 4.74 points, indicating that all users agree or strongly agree that our views’ design and the system’s interactive design are effective. S10 and S11 received about 82.61% strongly agree, demonstrating that Step View and Comparison View, designed specifically for T1-T5, are effective.
Usability (A4) To evaluate the usability of SingVisio, we design three related questions (S13-S15). Participants in the user study included those with over three years of experience in machine learning and signal processing, some new to these fields, and others with a purely musical background. Over half of the users strongly believed in the user-friendly interface and ease of learning of SingVisio (S13 and S15). However, the presence of strong disagreement in S13 and S15 indicates that our system still requires improvements to enhance its user-friendliness and ease of use. More than 78% users strongly recommended SingVisio to others who could benefit from its use (S14). This demonstrates that our system is user-friendly for diverse users, regardless of their background.
@@ 9 Conclusion
In this work, we introduce SingVisio, a visual analysis system designed to interactively explain the diffusion model for singing voice conversion. Specifically, SingVisio visually exhibits the step-wise generation process of diffusion models, illustrating the gradual denoising of the noisy spectrum, ultimately resulting in a clean spectrum that captures the target singer’s timbre. The system also supports pairwise comparisons between different conditions, such as content and melody in source audio, and timbre from the target audio, revealing the impact of these conditions on the diffusion generation process and converted results. Comparative and comprehensive evaluations demonstrate that SingVisio is effective in terms of system design, functionalities, explainability, and usability. It provides diverse users with fresh learning experiences and valuable insights into the diffusion model for singing voice conversion.
Appendix A Metrics Definition
  • 1.
  • 2.
  • 3.
  • 4.
  • 5.
@@ Appendix B Dataset
We use five datasets for our diffusion-based SVC model training: Opencpop [58], SVCC training data [49], VCTK [59], OpenSinger [60], and M4Singer [61]. In total, these datasets contain 83.1 hours of speech and 87.2 hours of singing data. The mapping between the singer name defined in the dataset and the singer ID displayed in the SingVisio system is listed in Table 1. The mapping between the song name defined in the dataset and song ID in the SingVisio system is listed in Table B.
@@
SVCC 30001 Song 1 Hey Jude, don’t make it bad.
30002 Song 2 Take a sad song and make it better.
30003 Song 3 Remember to let her into your heart.
10001 Song 4 Everything is fine.
10030 Song 5 Were you lying all the time?
10120 Song 6 Now, I need
10140 Song 7 Hey, I love you.
30005 Song 15 You know that its fool who plays it cool.
30006 Song 16 Na, na, na, na, na, na, na, na, na, na, na, na, hey, Jude.
30009 Song 17 When they all should let us be.
30016 Song 18 Let it be. Let it be. Let it be.
30022 Song 19 Take my breath away
30019 Song 20 Watching every motion In my foolish lover’s game
[…]
Fig. 7: The architecture of DiffWaveNetSVC [50, 19]. We select the Step, Step+Noise, and Step+Noise+Condition to project and visualize in SingVisio’s Project View (Section 5.4)
Appendix C Architecture of Diffusion-based SVC
The architecture of DiffWaveNetSVC is shown in Fig. 7.
Fig. 8: SingVisio User Study 1/7
Fig. 9: SingVisio User Study 2/7
Fig. 10: SingVisio User Study 3/7
Fig. 11: SingVisio User Study 4/7
Fig. 12: SingVisio User Study 5/7
Fig. 13: SingVisio User Study 6/7
Fig. 14: SingVisio User Study 7/7